factoring: quadratic equations and completing the square
a^6-4a^3-21=0

Answer by stanbon(75887)

a^6-4a^3-21=0
(a^3-7)(a^3+3) = 0
a = cube root of 7 or a = cube root of 3
============================================
Cheers,
Stan H.